OpenStudy (anonymous):

x=5t^2 y=t+4 find the graph of the curve

OpenStudy (anonymous):

re arrange x = 5t^2 for t t = sqroot ( x / 5 ) now sub into y = t + 4 , sub in first equation for t y = sqroot ( x / 5 ) + 4

OpenStudy (anonymous):

now graph the above function

OpenStudy (anonymous):

and you can see your curve

OpenStudy (anonymous):

is everything under the radical?

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

what do you take the square root of?

OpenStudy (anonymous):

\[y = \sqrt{(x / 5)} + 4\]

OpenStudy (kropot72):

\[y=\sqrt{\frac{x}{5}}+4\]

OpenStudy (anonymous):

ok awesome thanks